# Effect of Magnetic Field on the Formation of Radiatively Inefficient Accretion Flow around Black Holes

### v1

@@ III.3 Viscosity
Viscosity has been prescribed using the Shakura-Sunyaev viscosity prescription [4], where the viscous stress is considered as a perturbation of the local thermal pressure, the perturbation being specified by the dimensionless parameter $\alpha$. In this setup, only the $\tau_{r\phi}=\tau_{\phi r}$ of the viscous stress tensor is important since this is the component responsible for angular momentum transfer and hence accretion. This component of the viscous stress is given by
@@
$\eta$ is the dynamic viscosity coefficient. $\mu$ is related to the kinematic viscosity coefficient as $\mu=\rho\nu$. In the ’$\alpha$ disk’ model, $\nu$ is given by $\nu=\alpha H^{2}\Omega$, where $\Omega$ is the angular velocity and $H$ is the disk scale height. Thus,
@@
where a subscript of 00 indicates values at the torus center and $\alpha_{v}$ is a dimensionless number $<1$.
@@ III.6 Resistivity
The Shakura-Sunyaev prescription is used for resistivity as well. The resistivity $(\eta_{m})$ is given by
@@
Where $\nu_{m}$ is the magnetic diffusivity, which, according to the Shakura-Sunyaev parametrization, is given as
@@ VII Discussions & Conclusion
In this study, we investigated the numerical evolution of both magnetized and non-magnetized, radiatively inefficient accretion flows within a parametric framework to examine how magnetic fields influence the formation of thin disks. To explore this, we varied the strength of the seed magnetic field and the magnetic diffusivity and analyzed their impact on the flow characteristics. We see an increase in thin disk temperature with an increase in initial seed magnetic field strength. This increase in temperature with an increase in seed magnetic field strength can be attributed to a higher reconnection rate as a result of higher Alfven speeds. Our findings reveal a critical seed magnetic field value of $\sim$52 gauss at an accretion rate of $10^{-5}\ M_{\odot}/\text{year}$ and Shakura-Sunyaev magnetic diffusivity parameter ($\alpha_{m}$) of 0.03. Beyond this limit, the formation of a thin disk is not favourable, as indicated by our simulations and shown below. The above result can be attributed to the increased heating rate caused by the magnetic fields. To address the excess heating rate, we have introduced a new time scale, $\tau_{eq}$, which incorporates the time scales of heat sources involving viscosity, resistivity, magnetic reconnection and MRI (Eq. 32). We propose that for a thin disk to exist, this time scale must exceed the cooling time scale. By comparing the threshold values from our model with simulation results, we validate that our model accurately describes the behaviour of such accretion flows.
@@ VII.1 Estimation of the threshold value of plasma beta
We have estimated a threshold value for plasma beta, below which a geometrically thin accretion flow cannot be sustained. Since we have considered Shakura-Sunyaev parametrization for both resistivity and viscosity, both the timescales will be similar to each other (see Eq. 22 and Eq. 24). We propose that the timescale of reconnection must be equal to or lower than that of the resistive timescale to affect the nature of the flow. The cooling time scale would counter a time scale of higher order. Thus using Eq. 29, and Eq. 30, we propose:
@@
$\beta=\dfrac{K\rho^{\gamma-1}}{2\pi^{2}\eta^{2}}$ (35)
Equation 35 defines a threshold of plasma beta. If the material being transferred from a companion star onto a central object obeys a polytropic equation of state has a value of $\beta$ equal to or lower than the threshold provided by Eq. 35, it will only form a RIAF and the geometrically thin, optically thick disk will not be observed. From our simulations, the value of $\beta$ is $622$, while our model predicts a threshold beta value of $757$, which is of a similar order of magnitude. The slight difference in the value of $\beta$ can be attributed to numerical artefacts such as numerical diffusion. This is in accordance with observation where the source spends most of its time in the low-hard state rather than the high-soft state [45]. Alternatively, if the magnetic field gets amplified by the dynamo process due to an MRI, the excess heat generated in the inner region of the disk due to resistivity and reconnection at the midplane can cause a transition from a thin to thick accretion flow.

### v3

@@ III.3 Viscosity
The viscosity in our work has been prescribed using the viscous stress tensor. In this setup, only the $\tau_{r\phi}=\tau_{\phi r}$ of the viscous stress tensor is important since this is the component responsible for angular momentum transfer and hence accretion. Moreover, this is the most dominant component of magnetic stress in the magnetic stress tensor, which we have defined in Sec. III.6 [46]. This component of the viscous stress is given by
@@
$\eta$ is the dynamic viscosity coefficient. $\eta$ is related to the kinematic viscosity coefficient as $\eta=\rho\nu$.
The Shakura-Sunyaev $\alpha$ thin disk model specifies $\nu$ as $\nu=\alpha c_{s}H$ [4], where $c_{s}$ is the local sound speed. For a thin accretion disk, this is assumed to be constant, which is not true in general. Alternatively, the disk height can be expressed as $H(r)\approx c_{s}/\Omega(r)$ where $\Omega(r)$ is the angular velocity at $r$. This means, $\nu=\alpha H^{2}\Omega$ which is similar to the $\beta$ prescription provided in Eq. 1 of [47]. For a Keplerian orbit, $\Omega\propto r^{-3/2}$ and assuming a constant aspect ratio, the kinematic viscosity parameter $\nu$ varies as $r^{1/2}$ [9]. Thus, in this work, we have considered the following form of kinematic viscosity:
@@
where a subscript of 00 indicates values at the torus center and $\alpha_{v}$ is a dimensionless number $<1$ This means that $\nu_{v}\propto r^{1/2}\implies\eta\propto\rho r^{1/2}$ throughout the computational domain and is not constant.
@@ III.6 Resistivity
A prescription similar to the viscosity is used for resistivity as well. The resistivity $(\eta_{m})$ is given by
@@
Where $\nu_{m}$ is the magnetic diffusivity, which, according to the $\alpha$, is given as
@@ VII Discussions & Conclusion
In this study, we investigated the numerical evolution of both magnetized and non-magnetized, radiatively inefficient accretion flows within a parametric framework to examine how magnetic fields influence the formation of thin disks. To explore this, we varied the strength of the seed magnetic field and the magnetic diffusivity and analyzed their impact on the flow characteristics. We see an increase in thin disk temperature with an increase in initial seed magnetic field strength. This increase in temperature with an increase in seed magnetic field strength can be attributed to a higher reconnection rate as a result of higher Alfven speeds. Our findings reveal a critical seed magnetic field value of $\sim$52 gauss at an accretion rate of $10^{-5}\ M_{\odot}/\text{year}$ and $\alpha$ magnetic diffusivity parameter ($\alpha_{m}$) of 0.03. Beyond this limit, the formation of a thin disk is not favourable, as indicated by our simulations and shown below. The above result can be attributed to the increased heating rate caused by the magnetic fields. To address the excess heating rate, we have introduced a new time scale, $\tau_{eq}$, which incorporates the time scales of heat sources involving viscosity, resistivity, magnetic reconnection and MRI (Eq. 32). We propose that for a thin disk to exist, this time scale must exceed the cooling time scale. By comparing the threshold values from our model with simulation results, we validate that our model accurately describes the behaviour of such accretion flows.
@@ VII.1 Estimation of the threshold value of plasma beta
We have estimated a threshold value for plasma beta, below which a geometrically thin accretion flow cannot be sustained. Since we have considered $\alpha$ parametrization for both resistivity and viscosity, both the timescales will be similar to each other (see Eq. 22 and Eq. 24). We propose that the timescale of reconnection must be equal to or lower than that of the resistive timescale to affect the nature of the flow. The cooling time scale would counter a time scale of higher order. Thus using Eq. 29, and Eq. 30, we propose:
@@
$\beta_{th}=\dfrac{K\rho^{\gamma-1}}{2\pi^{2}\eta^{2}}$ (35)
Equation 35 provides a means to estimate the plasma beta beyond which a thin disk will be observed. If the material being transferred from a companion star onto a central object obeys a polytropic equation of state has a value of $\beta$ equal to or lower than the threshold provided by Eq. 35, it will only form a RIAF and the geometrically thin, optically thick disk will not be observed. This $\beta$ is a function of position and a local value of $\beta$ greater than this threshold can give rise to local collapses. However, this interpretation may be misleading. The correct way to view this is by considering the value of $\beta$ at a position far away from the compact object ($100R_{g}$ in our case). Since both resistivity, viscosity and pressure are varying (as a function of $r$ in our case), the values at that particular distance is to be considered. If the $\beta$ value is less than the threshold at this position, the evolution of the accretion flow will end up in a RIAF rather than forming a thin disk. In our case, ρ
(1/r)1/γ1
proportional-to
[…]
from Eq. 5 and $\eta\propto\rho r^{1/2}$. Using these conditions and Eq. 35, the $r$ dependence of $\beta$ in our case is given by
$\beta_{th}(r)\propto r^{\left[\dfrac{\gamma+1}{2(\gamma-3)}\right]}$ (36)
Eq. 36 shows the non-linear dependence of $\beta$ on the coordinate $r$. Thus the threshold is not a constant one but varies according to the viscosity coefficient and density which depends on $r$ in our case. From our simulations, the value of $\beta$ is $622$, while our model predicts a threshold beta value of $757$, which is of a similar order of magnitude. The slight difference in the value of $\beta$ can be attributed to numerical artifacts such as numerical diffusion. This is in accordance with observation where the source spends most of its time in the low-hard state rather than the high-soft state [45]. Alternatively, if the magnetic field gets amplified by the dynamo process due to an MRI, the excess heat generated in the inner region of the disk due to resistivity and reconnection at the midplane can cause a transition from a thin to thick accretion flow.